ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 102, 1997

Adams, T., N. Friedman and C. E. Silva Rank-one weak mixing for nonsingular transformations

269

Babson, E. K., L. J. Billera and C. S. Chan Neighborly cubical spheres and a cubical lower bound conjecture

297

Baruch, E. M. On the gamma factors attached to representations of U(2, 1) over a p-adic field

317

Bergman, G. M. Correction to “On the growth of algebras with bialgebra action” by George M. Bergman, Israel Journal of Mathematics, Vol. 96, 1996, pp. 63–96

347

Billera, L. J. See E. K. Babson

Bisztriczky, T. Ordinary (2m + 1)-polytopes  101
Blokh, A. and M. Misiurewicz Entropy of twist interval maps  61

Bump, D., S. Friedberg and M. Furusawa Explicit formulas for the Waldspurger and Bessel models

125

Chan, C. S. See E. K. Babson

Erdős, P., A. Hajnal and J. Pach On a metric generalization of Ramsey’s theorem

283

Ferenczi, V. A uniformly convex hereditarily indecomposable Banach space

199

Friedberg, S. See D. Bump

Friedman, N. See T. Adams

Furusawa, M. See D. Bump

Gelaki, S. Quantum groups of dimension pq^{2}  227

Glasner, E. A simple characterization of the set of μ-entropy pairs and applications

13

Gutiérrez, J. M. and J. G. Llavona Polynomially continuous operators

179

Hajnal, A. See P. Erdős

Llavona, J. G. See J. M. Gutiérrez

Lonke, Y. On zonoids whose polars are zonoids  1
Matoušek, J. On embedding expanders into ℓ_{p} spaces  189

Misiurewicz, M. See A. Blokh Murata, M. Semismall perturbations in the Martin theory for elliptic

equations

29

Pach, J. See P. Erdős Silva, C. E. See T. Adams